# Contact non-squeezing in lens spaces

### v1

@@ 6.1. Orderability and pseudo-spectral selectors
In particular for all non-empty open set $U$ in $M$ and $\psi\in\mathrm{Cont}_{0}(M,\xi)$
$\lceil C(U)\rceil_{T}=\lceil C(\psi(U))\rceil_{T},$
where $C(V):=\sup\{c(\widetilde{\varphi})\ |\ \widetilde{\varphi}\in\widetilde{\mathrm{Cont}_{0}^{\mathrm{c}}}(V,\xi)\}$ for all non-empty open sets $V\subset M$.
@@
6.2. Prequantization
Let $(W,\omega)$ be a closed symplectic manifold so that the natural morphism $H^{2}(W,\mathbb{Z})\to H^{2}(W,\mathbb{R})$ is without torsion, i.e. it is injective, $[\frac{\omega}{2\pi}]$ lies in its image and $[\omega],H_{2}(W,\mathbb{Z})\right>=2\pi\mathbb{Z}$. Then there exists a unique, up to $\mathbb{S}^{1}$-principal bundle isomorphism, $\mathbb{S}^{1}$-principal bundle $\pi_{1}:M_{1}\to W$ whose Euler class $eu(\pi_{1}:M_{1}\to W)=-[\frac{\omega}{2\pi}]$, where $\mathbb{S}^{1}=\{e^{i\theta}\ |\ \theta\in[0,2\pi]$}. Moreover $M_{1}$ carries a contact form $\alpha_{1}$ that satisfies $\pi_{1}^{*}\omega=\mathrm{d}\alpha_{1}$, the Reeb flow of which is Zoll of minimal period $2\pi$ and induces the $\mathbb{S}^{1}$ action on $M_{1}$. Finally, for any integer $k\geq 1$, seeing $\mathbb{Z}_{k}$ as the subgroup of $k$th roots of unity in $\mathbb{S}^{1}$ we have a bundle $\pi_{k}:M_{k}\to W$, where $M_{k}:=M_{1}/\mathbb{Z}_{k}$, with Euler class $-k[\frac{\omega}{2\pi}]$. This manifold $M_{k}$ carries a contact form $\alpha_{k}$ defined by $(\pi^{k}_{1})^{*}\alpha_{k}=\alpha_{1}$ where $\pi^{k}_{1}:M_{1}\to M_{k}$ is the covering map and it satisfies $\mathrm{d}\alpha_{k}=\pi_{k}^{*}\omega$. Moreover the Reeb flow of $\alpha_{k}$ is Zoll of minimal period $\frac{2\pi}{k}$ (see for instance the introduction of [34] and [19, Thm 7.2.4]). For all $k\in\mathbb{N}_{>0}$ we denote by $\xi_{k}:=\ker\alpha_{k}$ the contact distribution on $M_{k}$ and fix the coorientation given by the Reeb vector field.
@@ 6.2. Prequantization
Suppose there exists $k\in\mathbb{N}_{>0}$ such that $(M_{k},\xi_{k})$ is orderable. Then $(M_{Nk},\xi_{kN})$ is orderable for any $N\in\mathbb{N}_{>0}$. Moreover for any non-empty open set $U\subset W$
$C_{\alpha_{Nk}}(\pi_{Nk}^{-1}(U))\geq C_{\alpha_{k}}(\pi_{k}^{-1}(U)).$ (11)
@@
We leave the proof of the first statement to the reader. Let us thus prove the inequality (11). First note that $c_{+}^{\alpha_{Nk}}(\widetilde{\varphi})\geq c_{+}^{\alpha_{k}}(\mathcal{L}(\widetilde{\varphi}))$ for all $\widetilde{\varphi}\in\widetilde{\mathrm{Cont}_{0}}(M_{Nk},\xi_{Nk})$ where $\mathcal{L}:\widetilde{\mathrm{Cont}_{0}}(M_{Nk},\xi_{Nk})\to\widetilde{\mathrm{Cont}_{0}}(M_{k},\xi_{k})$ denotes the lifting map. Then remark that for all $\widetilde{\varphi}\in\widetilde{\mathrm{Cont}_{0}^{\mathrm{c}}}(\pi_{k}^{-1}(U),\xi_{k})$ there exists $\widetilde{\psi}\in\widetilde{\mathrm{Cont}_{0}^{\mathrm{c}}}(\pi_{Nk}^{-1}(U),\xi_{Nk})$ whose lift $\mathcal{L}(\widetilde{\psi})$ lies in $\widetilde{\mathrm{Cont}_{0}^{\mathrm{c}}}(\pi_{k}^{-1}(U),\xi_{k})$ and is such that $\widetilde{\varphi}\preceq\mathcal{L}(\widetilde{\psi})$. ∎
We have therefore the next Proposition.
@@
Let $\Psi:(\mathbb{B}^{2n}(R),\omega_{0})\hookrightarrow(W,\omega)$ be a symplectic embedding and suppose that $(M_{k},\ker\alpha_{k})$ is orderable for some integer $k$. If Conjecture 6.1 holds then there exists $N_{0}$ and $a_{0}\in(0,R]$ so that for all $N>N_{0}$ the open set $B_{Nk}(R)$ cannot be contactly squeezed inside $B_{Nk}(a)$ for all $a\in(0,a_{0}]$ where $B_{Nk}(r):=\pi_{Nk}^{-1}(\Psi(\mathbb{B}^{2n}(R))$ are seen as subset of the contact manifold $(M_{Nk},\ker\alpha_{Nk})$.
@@
Put $\delta:=C_{\alpha_{k}}(B_{k}(R))>0$. For $N_{0}\in\mathbb{N}_{>0}$ big enough $\delta>\frac{2\pi}{kN_{0}}$. Since, for any $N\in\mathbb{N}_{>0}$, $(M_{Nk},\xi_{Nk})$ is orderable and since we have assumed that Conjecture 6.1 holds it implies that one can apply the capacity-energy inequality of Section 4 and the computation in 5.2 to get that
$C_{\alpha_{Nk}}(B_{Nk}(a))\leq\pi a^{2}\leq\frac{\pi}{kN_{0}}$
for all $a\in(0,a_{0}]$, where $a_{0}:=\min\{\frac{R}{\sqrt{3}},(kN_{0})^{-\frac{1}{2}}\}$. In addition to that, thanks to Lemma 6.3
$\frac{2\pi}{kN_{0}}<C_{\alpha_{k}}(B_{k}(R))\leq C_{\alpha_{Nk}}(B_{Nk}(R)).$
Suppose by contradiction that there exists $\phi\in\mathrm{Cont}_{0}(M_{Nk},\xi_{k})$ so that $\phi(B_{Nk}(R))\subset B_{Nk}(a)$ for some integer $N>N_{0}$. Then by (10)
$C_{\alpha_{Nk}}(\phi(B_{Nk}(R)))\leq C_{\alpha_{Nk}}(B_{Nk}(a))\leq\frac{\pi}{kN_{0}}.$
Moreover, under the assumption that Conjecture 6.1 holds, one can use Lemma 6.2 to get that
$\lceil C_{\alpha_{Nk}}(\phi(B_{Nk}(R)))\rceil_{\frac{2\pi}{Nk}}=\lceil C_{\alpha_{Nk}}(B_{Nk}(R))\rceil_{\frac{2\pi}{Nk}}.$
We therefore get the following contradiction
$\frac{2\pi}{Nk}\left\lceil\frac{N}{N_{0}}\right\rceil=\left\lceil\frac{2\pi}{kN_{0}}\right\rceil_{\frac{2\pi}{Nk}}\leq\lceil C_{\alpha_{Nk}}(\phi(B_{Nk}(R)))\rceil_{\frac{2\pi}{Nk}}\leq\left\lceil\frac{\pi}{kN_{0}}\right\rceil_{\frac{2\pi}{Nk}}=\frac{2\pi}{Nk}\left\lceil\frac{1}{2}\cdot\frac{N}{N_{0}}\right\rceil.$

### v2

@@ 7.3. Order spectral selectors for strongly orderable prequantizations
If $(M_{1},\xi_{1})$ is a strongly orderable prequantization then for any $k\in\mathbb{N}_{>0}$
  1. (1)
    $(M_{k},\xi_{k})$ is strongly orderable
  2. (2)
    for any $\widetilde{\varphi}\in\widetilde{\mathrm{Cont}_{0}}(M_{k},\xi_{k})$ denoting by $\widetilde{\psi}\in\widetilde{\mathrm{Cont}_{0}}(M_{1},\xi_{1})$ its lift we have
    $C_{+}^{\alpha_{k}}(\widetilde{\varphi})\geq C_{+}^{\alpha_{1}}(\widetilde{\psi}).$
  3. (3)
    $\left\lceil C_{+}^{\alpha_{k}}(\widetilde{\psi}\widetilde{\varphi}\widetilde{\psi}^{-1})\right\rceil_{\frac{2\pi}{k}}=\left\lceil C_{+}^{\alpha_{k}}(\widetilde{\varphi})\right\rceil_{\frac{2\pi}{k}}$.
@@
In the following $(M_{k}\times M_{k}\times\mathbb{R},\Xi_{k}=\ker\beta_{k})$ denote the contact product of $(M_{k},\xi_{k}=\ker\alpha_{k})$ with itself and $\beta_{k}$ the product contact form associated to $\alpha_{k}$ for any $k\in\mathbb{N}_{>0}$ as in Section 1.2.
  1. (1)
    Suppose that $(M_{k},\xi_{k})$ is not strongly orderable. It means that there exists a positive contractible loop of Legendrians $(\Lambda_{t})\subset(M_{k}\times M_{k}\times\mathbb{R},\Xi_{k})$ based at $\Delta_{k}:=\{(x,x,0)\ |\ x\in M_{k}\}$. Consider $(\Lambda_{t}^{s})$ a homotopy of this loop with fixed endpoint to the constant loop based at $\Delta_{k}$, i.e. $\Lambda_{t}^{0}=\Lambda_{t}$ and $\Lambda_{t}^{1}=\Lambda_{1}^{s}=\Lambda_{0}^{s}=\Delta_{k}$ for all $t,\ s\in[0,1]$, and consider $(\Phi_{t}^{s})\subset\mathrm{Cont}_{0}(M_{k}\times M_{k}\times\mathbb{R},\Xi_{k})$ so that $\Phi_{t}^{s}(\Delta_{k})=\Lambda_{t}^{s}$ and $\Phi_{0}^{s}=\mathrm{id}$ for all $t,s\in[0,1]$ (for instance see [3, Lemma 2.1] for the construction of $(\Phi_{t}^{s})$). Seeing $M_{k}\times M_{k}\times\mathbb{R}$ as the quotient of $M_{1}\times M_{1}\times\mathbb{R}$ by the action by strict contactomorphisms of $\mathbb{Z}_{k}\times\mathbb{Z}_{k}$, i.e. $([j_{1}],[j_{2}])\cdot(x_{1},x_{2},\theta)=([j_{1}]\cdot x_{1},[j_{2}]\cdot x_{2},\theta)$, the group $\mathrm{Cont}_{0}^{\mathbb{Z}_{k}\times\mathbb{Z}_{k}}(M_{1}\times M_{1}\times\mathbb{R},\Xi_{1})$ is a covering space over $\mathrm{Cont}_{0}(M_{k}\times M_{k}\times\mathbb{R},\Xi_{k})$ and therefore there exists a unique $(\Psi_{t}^{s})\subset\mathrm{Cont}_{0}^{\mathbb{Z}_{k}\times\mathbb{Z}_{k}}(M_{1}\times M_{1}\times\Xi_{1})$ that lifts $(\Phi_{t}^{s})$, i.e. $\Pi_{k}\circ\Psi_{t}^{s}=\Phi_{t}^{s}\circ\Pi_{k}$ where $\Pi_{k}:M_{1}\times M_{1}\times\mathbb{R}\to M_{k}\times M_{k}\times\mathbb{R}$ denotes the natural projection, and so that $\Psi_{0}^{s}=\mathrm{id}$ for all $s,t\in[0,1]$. It is clear that $(\Psi_{t}^{1}(\Delta_{1}))$ is a contractible loop in $\mathcal{L}(\Delta_{1})$, and moreover since $\Pi_{k}^{*}\beta_{k}=\beta_{1}$ a direct computation shows that it is a positive loop which implies that $(M_{1},\xi_{1})$ is not strongly orderable.
  2. (2)
    Let $\widetilde{\varphi}\in\mathrm{Cont}_{0}(M_{k},\xi_{k})$ and its graph $\widetilde{\Phi}\in\mathrm{Cont}_{0}(M_{k}\times M_{k}\times\mathbb{R},\Xi_{k})$. Denoting by $\widetilde{\psi}$ the lift $\widetilde{\varphi}$ to $\mathrm{Cont}_{0}(M_{1},\xi_{1})$, it is easy to see that its graph $\widetilde{\Psi}\in\mathrm{Cont}_{0}(M_{1}\times M_{1},\times\mathbb{R},\Xi_{1})$ corresponds to the $\mathbb{Z}_{k}\times\mathbb{Z}_{k}$-equivariant lift of $\widetilde{\Phi}$. Consider any contact isotopy $(\eta_{t})\subset\mathrm{Cont}_{0}(M_{k}\times M_{k}\times\mathbb{R},\Xi_{k})$ starting at the identity such that $[(\eta_{t})]\cdot\Delta_{k}=\widetilde{\Phi}\cdot\Delta_{k}$, and $(\overline{\eta}_{t})\subset\mathrm{Cont}_{0}^{\mathbb{Z}_{k}\times\mathbb{Z}_{k}}(M_{1}\times M_{1}\times\mathbb{R},\Xi_{1})$ its corresponding $\mathbb{Z}_{k}\times\mathbb{Z}_{k}$-equivariant lift starting at the identity. We claim that $[(\overline{\eta_{t}})]\cdot\Delta_{1}=\widetilde{\Psi}\cdot\Delta_{1}$. Admitting this claim for the moment finishes the proof since $\underset{p\in\Delta_{1}}{\max}\beta_{1}(\frac{d}{dt}\overline{\eta}_{t}(p))=\underset{p\in\Delta_{k}}{\max}\beta_{k}(\frac{d}{dt}\eta_{t}(p))$. To prove the claim take any two contact isotopies $(\Phi_{t})$ and $(\Psi_{t})$ representing $\widetilde{\Phi}$ and $\widetilde{\Psi}$ respectively, i.e. $[(\Phi_{t})]=\widetilde{\Phi}$ and $[(\Psi_{t})]=\widetilde{\Psi}$. Then the Legendrian isotopy $(\Pi_{k}(\Psi_{t}^{-1}\circ\overline{\eta}_{t})\Lambda_{1})=(\Phi_{t}^{-1}\circ\eta_{t}(\Lambda_{k}))$ is a contractible loop based at $\Delta_{k}$ in $\mathcal{L}(\Delta_{k})$ since $[(\eta_{t})]\cdot\Delta_{k}=\widetilde{\Phi}\cdot\Delta_{k}$. Therefore the Legendrian isotopy $((\Psi_{t}^{-1}\circ\overline{\eta}_{t})\Lambda_{1})$ has to be itself a contractible loop in $\mathcal{L}(\Delta_{1})$ which concludes the proof of our claim.
  3. (3)
    It is a direct consequence of 2.2.
8. Computation of the contact capacity in strongly orderable prequantizations
Taking the same notations as in Section 7.3 above, suppose from now on that $(M_{1},\xi_{1}=\ker\alpha_{1})$ is a closed strongly orderable prequantization over $(W,\omega)$. Thanks to Lemma 7.5 the associated contact manifold $(M_{k},\xi_{k})$ is also strongly orderable, for all $k\in\mathbb{N}_{>0}$, and therefore one can define the spectral selectors $C_{+}^{\alpha_{k}}:\widetilde{\mathrm{Cont}_{0}}(M_{k},\xi_{k})\to\mathbb{R}$ of Section 7.2). Then, as in Section 4, for any $k\in\mathbb{N}_{>0}$999here $k$ is not necessarily greater than $1$ and any non-empty open set $U\subset M_{k}$ we denote by $\widetilde{\mathrm{Cont}_{0}^{\mathrm{c}}}(U,\xi_{k})$ the subgroup of $\widetilde{\mathrm{Cont}_{0}}(M_{k},\xi_{k})$ whose elements can be represented by contact isotopies $(\phi_{t})$ supported in $U$. We define similarly $C_{\alpha_{k}}:\{\text{non-empty open sets of }M_{k}\}\to\mathbb{R}_{\geq 0}\cup\{+\infty\}$ to be the map:
$U\mapsto C_{\alpha_{k}}(U):=\sup\{C^{\alpha_{k}}_{+}(\widetilde{\varphi})\ |\ \widetilde{\varphi}\in\widetilde{\mathrm{Cont}_{0}^{\mathrm{c}}}(U,\xi_{k})\}.$
Recall that $\pi_{k}:M_{k}\to W$ denotes the natural projection for all $k\in\mathbb{N}_{>0}$. As for lens spaces we have the following:
Theorem 8.1.
For all non-empty open sets $U,V\subset M_{k}$ and $\phi\in\mathrm{Cont}_{0}(M_{k},\xi_{k})$
  1. (1)
    (monotonicity) $C_{\alpha_{k}}(U)\leq C_{\alpha_{k}}(V)$ if $U\subset V$
  2. (2)
    (invariance by conjugation) $\lceil C_{\alpha_{k}}(U)\rceil_{\frac{2\pi}{k}}=\lceil C_{\alpha_{k}}(\phi(U))\rceil_{\frac{2\pi}{k}}$
  1. (3)
    (lower bound) if $\Psi:(\mathbb{B}^{2n}(R),\omega_{0})\hookrightarrow(W,\omega)$ is a symplectic embedding for some $R\in(0,\sqrt{2}]$ then $C_{\alpha_{k}}(B_{k}(r))\geq\pi r^{2}$ for all $r\in(0,R]$ where $B_{k}(r):=\pi_{k}^{-1}(\Psi(\mathbb{B}^{2n}(r)))$,
  2. (4)
    (upper bound) if $\Psi:(\mathbb{B}^{2n}(\lambda R),\omega_{0})\hookrightarrow(W,\omega)$ is a symplectic embedding for some $R>0$ and $\lambda>\sqrt{3}$ then $C_{\alpha_{k}}(B_{k}(r))\leq\pi r^{2}$ for all $r\in(0,R]$ where $B_{k}(r):=\pi_{k}^{-1}(\Psi(\mathbb{B}^{2n}(r)))$.
To prove the monotonicity and the invariance by conjugation properties one can argue exactly as in Section 4. Also, note that as in Section 4, one deduces immediatly Theorem 1.3 from Theorem 8.1. It thus remains to prove the last two estimates and the proof will mimick the proof we have given in Section 6 with minor modifications.
8.1. Computation of the lower bound
Let us first prove the lower bound that we restate below.
Proposition 8.2.
If $\Psi:(\mathbb{B}^{2n}(R),\omega_{0})\hookrightarrow(W,\omega)$ is a symplectic embedding for $R\in(0,\sqrt{2}]$ then $C_{\alpha_{k}}(B_{k}(r))\geq\pi r^{2}$ for all $r\in(0,R]$ where $B_{k}(r):=\pi_{k}^{-1}(\Psi(\mathbb{B}^{2n}(r)))$.
Before proving Proposition 8.2 let us state the following Lemma.
@@ 8.1. Computation of the lower bound
Let $U$ be an open subset of $(W,\omega)$ then for any $k\in\mathbb{N}_{>0}$
$C_{\alpha_{k}}(\pi_{k}^{-1}(U))\geq C_{\alpha_{1}}(\pi_{1}^{-1}(U)).$
@@
For all $\widetilde{\varphi}_{1}\in\widetilde{\mathrm{Cont}_{0}^{\mathrm{c}}}(\pi_{1}^{-1}(U))$ one can easily construct $\widetilde{\psi}_{k}\in\widetilde{\mathrm{Cont}_{0}^{\mathrm{c}}}(\pi_{k}^{-1}(U))$ so that the lift of $\widetilde{\psi}_{k}$ to $\widetilde{\mathrm{Cont}_{0}^{\mathrm{c}}}(M_{1},\xi_{1})$ that we denote by $\widetilde{\psi}_{1}$ satisfies $\widetilde{\varphi}_{1}\preceq\widetilde{\psi}_{1}$ and $\widetilde{\psi}_{1}\in\widetilde{\mathrm{Cont}_{0}^{\mathrm{c}}}(\pi_{1}^{-1}(U))$. Moreover thanks to Lemma 7.5 and the monotonicity of $C_{+}^{\alpha_{1}}$ we have
$C_{+}^{\alpha_{k}}(\widetilde{\psi}_{k})\geq C_{+}^{\alpha_{1}}(\widetilde{\psi}_{1})\geq C_{+}^{\alpha_{1}}(\widetilde{\varphi}_{1})$
which concludes the proof. ∎
Proof of Proposition 8.2.
The first 4 steps of the proof will follow the same steps as the proof of Proposition 6.2.
Step 1 : For any $r\in(0,R]$ and $0<\varepsilon<\pi r^{2}$, consider the same function $f:\mathbb{R}_{\geq 0}\to\mathbb{R}_{\geq 0}$ as in the first step of the proof of Proposition 6.2 and consider the same map $H:\mathbb{B}^{2n}(r)\to\mathbb{R},\ x\mapsto f(\|x\|^{2})$.
Step 2 : Use the symplectic embedding $\Psi$ to import the construction to $(W,\omega)$, i.e.
$G:W\to\mathbb{R}\quad\quad x\mapsto\left\{\begin{array}[]{ll}H(\Psi^{-1}(x))&\text{ if }x\in\Psi(\mathbb{B}^{2n}(r))\\ 0&\text{ otherwise}\end{array}.\right.$
Step 3 : Import the construction to $(M_{1},\xi_{1})$ by considering the $\alpha_{1}$-Hamiltonian function
$h:M_{1}\to\mathbb{R},\quad\quad\quad x\mapsto G(\pi_{1}(x)).$
Therefore the same arguments as in the proof of Proposition 6.3 shows that for all $T\in[0,1]$
$\mathrm{Spec}^{\alpha_{1}}(\phi_{h}^{T})=\left\{T(\pi r^{2}-\varepsilon)+2\pi m\ |\ m\in\mathbb{Z}\right\}\bigcup\left\{2\pi m\ |\ m\in\mathbb{Z}\right\},$
where $\phi_{h}^{T}\in\mathrm{Cont}_{0}(M_{1},\xi_{1})$.
Step 4 : Therefore arguing as for Proposition 6.3 we get that $c(\widetilde{\varphi}_{T})=T(\pi r^{2}-\varepsilon)$ for all $T\in[0,\delta]$ where $\delta=\min\{1,\frac{2\pi}{\pi r^{2}-\varepsilon}\}$ and $\widetilde{\varphi}_{T}:=[(\phi_{h}^{tT})]$. Since by assumption $R\leq\sqrt{2}$ we deduce that $\delta=1$ and thus $c(\widetilde{\varphi}_{1})=\pi r^{2}-\varepsilon$ which implies that $C_{\alpha_{1}}(B_{1}(r))\geq\pi r^{2}-\varepsilon$. Taking the limit when $\varepsilon$ tends to 00 yields $C_{\alpha_{1}}(B_{1}(r))\geq\pi r^{2}$.
Step 5 : It remains now to apply Lemma 8.3 to deduce that
$C_{\alpha_{k}}(B_{k}(r))\geq C_{\alpha_{1}}(B_{1}(r))\geq\pi r^{2}.$
8.2. Computation of the upper bound
To finish the proof of Theorem 8.1 it thus remains to prove the upper bound that we restate below.
@@ 8.2. Computation of the upper bound
If $\Psi:(\mathbb{B}^{2n}(\lambda R),\omega_{0})\hookrightarrow(W,\omega)$ is a symplectic embedding for some $R>0$ and $\lambda>\sqrt{3}$ then $C_{\alpha_{k}}(B_{k}(r))\leq\pi r^{2}$ for all $r\in(0,R]$ where $B_{k}(r):=\pi_{k}^{-1}(\Psi(\mathbb{B}^{2n}(r)))$.
This proposition will be an adaptation of the proof of Proposition 6.3 together with the following Lemma which is an analogue of Theorem 5.3. First let us remark that
[…]
Lemma 8.5.
Let $U$ be an open subset of $M_{k}$ and $\widetilde{\psi}\in\widetilde{\mathrm{Cont}_{0}}^{\mathbb{T}^{1}}(M_{k},\xi_{k})$. If $\Pi(\widetilde{\psi})$ $\alpha_{k}$-displaces $U$ then $C_{\alpha_{k}}(U)\leq C_{+}^{\alpha_{k}}(\widetilde{\psi}^{-1})+C_{+}^{\alpha_{k}}(\widetilde{\psi}).$
@@
Let $\widetilde{\varphi}\in\widetilde{\mathrm{Cont}_{0}^{\mathrm{c}}}(U,\xi_{k})$. Then thanks to Proposition 5.5 $C_{+}^{\alpha_{k}}(\widetilde{\psi}\widetilde{\varphi})=C_{+}^{\alpha_{k}}(\widetilde{\psi})$. Moreover since $\widetilde{\psi}\in\widetilde{\mathrm{Cont}_{0}}^{\mathbb{T}^{1}}(M_{k},\xi_{k})$ it implies that $\widetilde{\psi}^{-1}\in\widetilde{\mathrm{Cont}_{0}}^{\mathbb{T}^{1}}(M_{k},\xi_{k})$. Hence Corollary 7.3 together with (14) ensure that
$C_{+}^{\alpha_{k}}(\widetilde{\varphi})=C_{+}^{\alpha_{k}}\left(\widetilde{\psi}^{-1}\cdot(\widetilde{\psi}\widetilde{\varphi})\right)\leq C_{+}^{\alpha_{k}}(\widetilde{\psi}^{-1})+C_{+}^{\alpha_{k}}(\widetilde{\psi}\widetilde{\varphi})=C_{+}^{\alpha_{k}}(\widetilde{\psi}^{-1})+C_{+}^{\alpha_{k}}(\widetilde{\psi})$
which concludes the proof. ∎
Proof of Proposition 8.4.
Let $r\in(0,R]$, $\delta>0$ be small enough so that $\sqrt{3+2\delta}<\lambda$.
Step 1 : Consider the Hamiltonian function $\widetilde{K}$ supported in $\mathbb{B}^{2n}(\lambda R)$ constructed in the proof of Proposition 6.3
  1. (1)
    $\widetilde{K}$ take values in $[0,r^{2}(\pi+\delta)]$
  2. (2)
    its time $1$ Hamiltonian flow displaces $B_{k}(r)$.
Step 2 : Import the construction to $(W,\omega)$ using $\Psi$, i.e.
$G:W\to\mathbb{R}\quad\quad x\mapsto\left\{\begin{array}[]{ll}H(\Psi^{-1}(x))&\text{ if }x\in\Psi(\mathbb{B}^{2n}(r))\\ 0&\text{ otherwise}\end{array}.\right.$
Step 3 : Import the construction to $(M_{k},\xi_{k})$ using Lemma 6.1. More precisely the $\alpha_{k}$-Hamiltonian function $h:M_{k}\to\mathbb{R}_{\geq 0}$, $x\mapsto G(\pi_{k}(x))$ generates a flow of $\mathbb{T}^{1}$-equivariant contactomorphisms $(\phi_{h}^{t})\subset\mathrm{Cont}_{0}(M_{k},\xi_{k})$. Moreover
  1. (1)
    $h\geq 0$ and $\max h\leq r^{2}(\pi+\delta)$
  2. (2)
Step 4 : Hence thanks to Lemma 8.5
$C_{\alpha_{k}}(B_{k}(r))\leq C_{+}^{\alpha_{k}}(\widetilde{\phi}^{-1})+C_{+}^{\alpha_{k}}(\widetilde{\phi})$
where $\widetilde{\phi}:=[(\phi_{h}^{t})]\in\widetilde{\mathrm{Cont}_{0}}^{\mathbb{T}^{1}}(M_{k},\xi_{k})$. Finally, since the $\alpha_{k}$-Hamiltonian function $M_{k}\to\mathbb{R}_{\leq 0},\ x\mapsto-h(x)$, generates the contact isotopy $(\phi_{h}^{t})^{-1}$ which represents $\widetilde{\phi}^{-1}$, Proposition 7.2 implies that
$C_{+}^{\alpha_{k}}(\widetilde{\phi}^{-1})+C_{+}^{\alpha_{k}}(\widetilde{\phi})\leq\int_{0}^{1}\max h\mathrm{d}t-\int_{0}^{1}\min h\mathrm{d}t\leq r^{2}(\pi+\delta).$
Letting $\delta$ goes to 00 finishes the proof.